A chemical company makes two brands of antifreeze. The first brand is 30% pure antifreeze, and the second brand is 55% pure antifreeze. In order to obtain 120 gallons of a mixture that contains 40% pure antifreeze,how many gallons of each brand of antifreeze must be used. Please help me solve this equation.
Answer by Cromlix(4381)

Hi there,
Amount.....Percentage....Total
solution 1...x .........30 ....... 30%x
solution 2.(120 - x)....55 ....... 55(120 - x)
solution 3...120........40........ 120(40)
30%x + 55%(120 - x) = 40%(120)
0.3x + 0.55(120 - x) = 0.4(120)
0.3x + 66 - 0.55x = 48
-0.25x + 66 = 48
-0.25x = 48 - 66
-0.25x = - 18
x = 72
Solution 1 = 72 gallons
Solution 2 = 48 gallons
Proof:
72 galls of 30% + 48 gallons of 55% = 120 galls of 40%
0.3(72) + 0.55(48) = 0.4(120)
21.6 + 26.4 = 48
Hope this helps :-)